# On Generating Monolithic and Model Reconciling Explanations in Probabilistic Scenarios

### v1

@@ 1 Introduction
Parallel to advancements in ML, the automated planning community has adopted a focused approach to generating explanations for plans produced by AI (planning) agents, which led to the inception of explainable AI planning (XAIP) (?). Predominantly, XAIP research focuses on the explanation generation problem, which involves identifying explanations for plans that, when conveyed to human users, help them understand and accept the agent’s proposed actions (?, ?). A noteworthy direction within this space is the model reconciliation problem (MRP) (?, ?, ?, ?), aimed at aligning a user’s model with that of an AI agent through the provision of explanations, especially when discrepancies in their understanding of a planning problem lead to confusion or misinterpretation of the agent’s decisions. We will refer to these explanations as model reconciling explanations. In contrast, we refer to explanations that do not account for the user’s model as monolithic explanations.
@@
Although MRP tackles essential facets of explainability, it often operates under the presumption that the AI agent has a deterministic grasp of the human model – a scenario that may not always align with the complexities of real-world interactions characterized by uncertainty about human knowledge. Indeed, this gap highlights a general challenge in explanation generation from AI agents: Their operation within realms of incomplete information and probabilistic decision-making models. Traditional explanation methods, which rely on deterministic knowledge, falter under these conditions, unable to represent the uncertain nature of the agent’s knowledge adequately.
@@
To bridge this gap in explainability, the first part of the paper introduces an approach that integrates uncertainty into the explanation generation framework. In particular, using (propositional) probabilistic logic as our underlying mechanism for modeling uncertainty, we propose a framework, where given a belief base $\mathcal{B}$, which is a weighted knowledge base, and an explanandum $\varphi$, the goal is to find a probabilistic monolithic explanation that increases the probability of the explanandum, that is, the probability that the explanandum is true. We introduce the concepts of explanatory gain and explanatory power, which serve as metrics for evaluating the effectiveness of probabilistic monolithic explanations. Furthermore, we extend our framework from logic-based model reconciliation problems (L-MRPs) (?, ?, ?), focusing on scenarios beset with uncertainty in the agent’s perception of the human model. Specifically, given a knowledge base $\mathrm{KB}_{\alpha}$ of an agent, an explanandum $\varphi$ entailed by $\mathrm{KB}_{\alpha}$, and a human belief base $\mathcal{B}_{h}$, the goal is to find a probabilistic model reconciling explanation that not only increases the explanandum’s probability for $\mathcal{B}_{h}$ but also decreases the probability of conflicts between the explanation and $\mathcal{B}_{h}$.
@@
In the second part of the paper, we describe algorithms for computing both probabilistic monolithic and model reconciling explanations, leveraging the conceptual duality between minimal correction sets (MCSes) and minimal unsatisfiable sets (MUSes). These algorithms, adapted from our prior work in ? (?), integrate a weighted maximum satisfiability procedure for computing probabilities. Our experimental evaluation across various benchmarks underscores the practicality and applicability of our algorithms, demonstrating their capability to generate probabilistic explanations within a propositional logic framework.
@@
    We propose a novel framework for generating probabilistic monolithic explanations given a belief base and an explanandum. Central to our framework are the concepts of explanatory gain and explanatory power, metrics designed to quantitatively assess the quality and effectiveness of probabilistic explanations.
@@
    We present a probabilistic framework for L-MRP to tackle scenarios characterized by uncertainty in the agent’s understanding of the human model.
@@
    We describe algorithms for computing probabilistic monolithic and model reconciling explanations, using the duality of MCSes and MUSes together with a weighted maximum satisfiability process.
@@
    Through a series of benchmark evaluations, we demonstrate the efficacy of our proposed algorithms in generating probabilistic explanations.
The paper is structured as follows. In Section 2, we provide the necessary background knowledge. In Section 3, we describe our explanation generation framework for monolithic and model reconciliation probabilistic explanations. In Section 4, we present algorithms for computing classical and probabilistic explanations, and experimentally evaluate them on a set of benchmarks in Section 5. Finally, in Section 6, we discuss related work, and conclude the paper in Section 7.
@@ 2.1 Propositional Logic
Consider the knowledge base $\mathrm{KB}=\{p,\lnot p\vee q,\lnot p\vee r\}$ build up from $\mathcal{V}=\{p,q,r\}$. Notice that $\mathrm{KB}\models q$. Then, $\epsilon=\{p,\lnot p\vee q\}$ is a monolithic explanation for $q$ from $\mathrm{KB}$.
Note that in this paper we do not consider formulae $\epsilon\equiv\varphi$ as monolithic explanations. These trivial explanations, which are of the form “why $\varphi$, because $\varphi$, are uninformative pertaining the explanandum.
@@ 2.1.1 Duality of Minimal Unsatisfiable and Minimal Corrections Sets
Partial MUSes in an inconsistent knowledge base $\mathrm{KB}$ appear when a subset of formulae is set as hard, that is, formulae that must always be satisfied in a solution. Conversely, soft formulae may not always be satisfied. Given a formula $\varphi$, we will write $\varphi^{*}$ with $*\in\{s,h\}$ to denote it as soft and hard, respectively.
MUSes and MCSes are related by the concept of minimal hitting set:
Definition 5 (Minimal Hitting Set).
Given a collection $\Gamma$ of sets from a universe $U$, a hitting set for $\Gamma$ is a set $H\subseteq U$ such that $\forall S\in\Gamma$, $H\cap S\neq\emptyset$ and $\nexists H^{\prime}\subset H$ such that $H^{\prime}\cap S\neq\emptyset$.
The relationship between MUSes and MCSes is discussed by ? (?) and ? (?), and it was firstly presented by ? (?), where MUSes and MCSes are referred to as (minimal) conflicts and diagnoses, respectively.
@@
Let $\mathrm{KB}=\{p,\lnot p\vee q,\lnot p\vee r\}$ and $\epsilon=\{p,\lnot p\vee q\}$ from Example 1. Notice how =$\mathcal{M}=\{p,\lnot p\vee q,\lnot q\}$ is an MUS of $\mathrm{KB}\cup\{\lnot q\}$. Then, it is easy to see that $\epsilon$ is a partial MUS of $\mathcal{M}$.
@@ 2.2 Modeling Uncertainty in Propositional Logic
Let $\Omega$ be the set of possible worlds and $P$ a probability distribution over $\Omega$. The degree of belief in a formula $\varphi\in\mathcal{L}$ is $P(\varphi)=\underset{\omega\models\varphi}{{{\sum}}}P(\omega)$.
@@
=$\mathcal{B}=\{(\phi_{1},w_{1}),\ldots,(\phi_{n},w_{n})\}$ (1)
where each formula $\phi_{i}\in\mathcal{L}$ is associated with a weight $w_{i}\in\mathbb{R}^{+}$. [^2] We assume, without loss of generality, that all weights are non-negative because a formula with a negative weight $w$ can be replaced by its negation with weight $-w$.
Intuitively, the weights serve as meta-information and reflect the certainty about the truth of the corresponding formulae – the higher the weight, the more certain the formula is. In that sense, formulae with higher weights are prioritized for satisfaction, effectively capturing the certainty of the particular formulae. This mechanism is especially useful for handling inconsistency and non-monotonic reasoning patterns, thus capturing a broader spectrum of problems. [^3] For example, the notion of inconsistency is relaxed as follows: Given two inconsistent formulae $\phi$ and $\lnot\phi$, if $P(\phi)=0.9$, then from the axioms of probability we have that $P(\lnot\phi)=0.1$. This then means that the worlds where $\lnot\phi$ is true are more unlikely than the worlds where $\phi$ is true, but not impossible. Further, we will denote with $\mathcal{B}^{\downarrow w}$ the classical projection of $\mathcal{B}$, that is, $\mathcal{B}^{\downarrow w}=\{\phi_{i}\>|\>(\phi_{i},w_{i})\in\mathcal{B}\}$.
@@
[…]
@@
[…]
@@ 2.3 The Model Reconciliation Problem
Now, building upon the MRP foundation, ? (?) introduced its logic-based variant (L-MRP), where the models of the agent and the human user are represented as logical knowledge bases. As a model reconciling explanation must take into account both the knowledge base $\mathrm{KB}_{\alpha}$ of the agent providing an explanation as well as the knowledge base $\mathrm{KB}_{h}$ of the human receiving the explanation, it is defined slightly differently compared to monolithic explanations defined by Definition 1:
@@
Let $\mathrm{KB}_{\alpha}=\{a,\lnot a\vee b,\lnot a\vee c\}$ and $\mathrm{KB}_{h}=\{\lnot a,\lnot a\vee b\}$ be the knowledge bases of an agent and a human user, respectively, where $\mathrm{KB}_{\alpha}\models b$ and $\mathrm{KB}_{h}\not\models b$. A model reconciling explanation is then =$\mathcal{E}=\langle\{a\},\{\lnot a\}\rangle$, where $(\mathrm{KB}_{h}\cup\{a\})\setminus\{\lnot a\}=\{a,\lnot a\vee b\}\models b$.

### v2

@@ 1 Introduction
Parallel to advancements in ML, the automated planning community has adopted a focused approach to generating explanations for plans produced by AI (planning) agents, which led to the inception of explainable AI planning (XAIP) (?). Predominantly, XAIP research focuses on the explanation generation problem, which involves identifying explanations for plans that, when conveyed to human users, help them understand and accept the agent’s proposed actions (?, ?). A noteworthy direction within this space is the model reconciliation problem (MRP) (?, ?, ?, ?, ?), aimed at aligning a user’s model with that of an AI agent through the provision of explanations, especially when discrepancies in their understanding of a planning problem lead to confusion or misinterpretation of the agent’s decisions. We will refer to these explanations as model reconciling explanations. In contrast, we refer to explanations that do not account for the user’s model as monolithic explanations.
@@
Although MRP tackles essential facets of explainability, it often operates under the presumption that the AI agent has a deterministic grasp of the human model—a scenario that may not always align with the complexities of real-world interactions characterized by uncertainty about human knowledge. Indeed, this gap highlights a general challenge in explanation generation from AI agents: Their operation within realms of incomplete information and probabilistic decision-making models. Traditional explanation methods, which rely on deterministic knowledge, falter under these conditions, unable to represent the uncertain nature of the agent’s knowledge adequately.
@@
To bridge this gap in explainability, the first part of the paper presents a framework for explanation generation under uncertainty. In particular, we use (propositional) probabilistic logic as our underlying mechanism for modeling uncertainty, and introduce the notions of probabilistic monolithic explanation and probabilistic model reconciling explanation. For the former, given an agent’s belief base $\mathcal{B}_{\alpha}$ (a weighted knowledge base [^1] A weighted knowledge base, essentially, induces a probability distribution over the variables of the agent’s language (?).), and an explanandum $\varphi$ (a formula to be explained), the goal is to find a probabilistic monolithic explanation such that the probability of the explanandum being true is increased. Moreover, we extend our work on logic-based model reconciliation problems (L-MRPs) (?, ?, ?) to scenarios in which the human user’s model is uncertain. Specifically, given a knowledge base $\mathrm{KB}_{\alpha}$ of an agent, an explanandum $\varphi$ entailed by $\mathrm{KB}_{\alpha}$, and a human belief base $\mathcal{B}_{h}$, the goal is to find a probabilistic model reconciling explanation such that the explanandum’s probability is increased while the probability of conflicts between the explanation and $\mathcal{B}_{h}$ is decreased. To measure the quality of such explanations, we define the concepts of explanatory gain and explanatory power, aimed at quantifying the effectiveness of these explanations with respect to the explanandum.
@@
In the second part of the paper, we describe algorithms for computing both types of explanations, where we leverage the duality between minimal correction sets (MCSes) and minimal unsatisfiable sets (MUSes). These algorithms, adapted from our prior work (?), integrate a weighted maximum satisfiability procedure for computing probabilities. Our experimental evaluation across various benchmarks highlights the practicality and applicability of our algorithms, and demonstrate their capabilities to generate probabilistic explanations within a propositional logic framework.
@@
    We propose a novel framework for generating probabilistic monolithic explanations and probabilistic model reconciling explanations. Central to our framework are the concepts of explanatory gain and explanatory power, metrics designed to quantitatively assess the quality and effectiveness of probabilistic explanations.
@@
    We describe algorithms for computing both types of probabilistic explanations, using the duality of MCSes and MUSes together with a weighted maximum satisfiability process. Through a series of benchmark evaluations, we demonstrate the efficacy of our proposed algorithms in generating explanations.
The paper is structured as follows. In Section 2, we motivate the use and applicability of logic in explainability, and discuss possible sources of uncertainty in explanation generation. In Section 3, we provide the background knowledge needed, and in Section 4 we describe a motivating application that serves as a running example. In Section 5, we present our explanation generation framework for monolithic and model reconciliation probabilistic explanations. We present algorithms for computing explanations in Section 6, and experimentally evaluate them on a set of benchmarks in Section 7. We then discuss related work in Section 8, address the assumptions, limitations and future extensions of our framework in Section 9, and finally conclude the paper in Section 10.
2 A Logic-based Perspective on Explainable Decision-Making
In this work, we take the following perspective: Logic-based frameworks can serve as an explanatory representational layer for AI systems, enabling the generation of rigorous and flexible explanations across diverse problem domains by capturing the system’s decisions in a formal logical language that supports inference and reasoning.
According to this perspective, formal logic provides a robust foundation for creating explanatory mechanisms by serving as an intermediate representational layer between AI systems and explanation generation processes. Some of the reasons that justify this perspective are as follows:
@@ 2 A Logic-based Perspective on Explainable Decision-Making
    Structured Semantics: Logic has well-defined compositional semantic functions that compute the meaning of a compound as a function of its constituents’ meanings. This composition is invertible, allowing us to trace back from conclusions to premises—essentially performing inference in reverse. This property is fundamental to explanation, as it enables us to identify precisely which components contributed to a particular decision or output.
@@
    Expressivity & Scrutability: Logic can represent complex knowledge and reasoning processes in a form that can be traced and scrutinized. Depending on the formal language of choice, it allows for encoding rich, relational information about the world, including classes, hierarchies, causal relationships, and quantified statements. For instance, in planning domains, logical representations can express not only states and actions but also the causal links between them, making it possible to explain why certain actions were selected in a plan. Moreover, logic-based systems enable us to examine their internal properties, both through internal verification techniques and external dialogues. This scrutability is crucial for building trust in AI systems, as it allows for thorough validation of the reasoning process, ensuring that explanations are not just plausible but provably correct within the system’s logical framework.
  • $\bullet$
    Adaptability: Logic-based systems can be extended with new knowledge and can integrate with various AI paradigms. This means that explanatory frameworks built on logic can evolve over time, incorporating new concepts, rules, or meta-level reasoning principles as needed. This adaptability is important for explanation systems that must operate across diverse domains or interact with different types of users.
  • $\bullet$
    Uncertainty Representation: While classical logic is deterministic, probabilistic extensions (which we use in this work) allow for representing and reasoning about uncertainty. This is important in real-world scenarios where AI systems must operate with incomplete information or where multiple interpretations of data are possible. Probabilistic logic enables explanations that convey not just what the system “believes”, but also its degree of confidence in those beliefs.
For more about logic and its suitability in explainability as well as the current AI landscape, please see some good arguments presented by ? (?), ? (?), and ? (?).
Figure 1: The Logic-based Explainability Layer. Given a decision $d$ from an AI model, a monolithic explanation is generated with respect to the AI system’s knowledge base ($\mathrm{KB}_{\alpha}$), while a model reconciling explanation is generated with respect to the AI system’s ($\mathrm{KB}_{\alpha}$) as well as the human user’s knowledge base ($\mathrm{KB}_{h}$).
Now, as illustrated in Figure 1, our approach positions logic as an intermediate explanatory layer between AI systems and human users. The logic-based explainability layer functions as an abstraction mechanism that captures the essential reasoning behind an AI system’s decisions in a formal logical representation, regardless of the underlying implementation details of the AI system itself. This layer serves two critical functions. First, it translates the internal decision processes of potentially opaque AI systems into a logical formalism that can be systematically analyzed. Second, it enables the generation of explanations through formal reasoning over this logical representation.
We will distinguish between two notions of explanation: monolithic explanations, which are generated with respect to the AI system’s knowledge base ($\mathrm{KB}_{\alpha}$) alone, focusing on providing a self-contained justification for the system’s decision; and model reconciling explanations, which are generated with respect to both the AI system’s knowledge base ($\mathrm{KB}_{\alpha}$) and the human user’s knowledge base ($\mathrm{KB}_{h}$), explicitly addressing the discrepancies between the two models that may lead to confusion or misunderstanding. The explainability layer thus provides a unified framework for generating different types of explanations while maintaining a clear separation between the underlying AI system and the explanation mechanism.
The effectiveness of the logic-based explainability framework depends crucially on the ability to encode the problem domains of interest in logical formalisms. In general, it can be used for problems that can be represented in a logical language for which satisfiability of sets is feasible. Fortunately, many problem domains admit logical encodings, such as planning problems (?, ?, ?), scheduling problems (?, ?, ?), argumentation problems (?, ?, ?), as well as machine learning problems (?, ?, ?). Indeed, logic-based explanation generation approaches have been successfully employed in numerous applications (?, ?, ?, ?, ?).
2.1 The Source of Uncertainty in Explanations
Traditional logic-based approaches have typically operated within deterministic settings. They assume that both the AI system and the human user have definite, certain knowledge about the domain. This assumption, however, fails to capture the reality of many real-world scenarios where uncertainty plays a crucial role. Succinctly,
  • $\bullet$
    Incomplete Information: AI systems often operate with partial information about the environment, meaning they operate with probabilistic rather than deterministic knowledge.
  • $\bullet$
    Ambiguous Evidence: In domains like medical diagnosis or law, the available evidence may support multiple hypotheses to varying degrees, requiring probabilistic reasoning.
  • $\bullet$
    Uncertain Human Models: Human users may hold beliefs with varying degrees of confidence, and AI systems may be uncertain about what humans know or believe.
  • $\bullet$
    Probabilistic Effects: In more realistic, dynamic environments, actions may have probabilistic outcomes, requiring explanations that account for these uncertainties.
Motivating Example. To ground our discussion, consider a service robot operating in an office building with some uncertain information about its environment. The robot must navigate from its current location to deliver a package to a specific office, choosing among multiple possible routes. In this scenario, uncertainty may arise due to several reasons, such as environmental uncertainty, e.g., the robot has limited information about the current state of different corridors (e.g., crowded or clear), outcome uncertainty, e.g., even with a chosen path, the robot may not be able to perfectly predict travel times due to potential obstructions or changing conditions, and uncertainty about the beliefs of humans, i.e., when explaining its decisions to building staff, the robot may not know exactly what they believe about current building conditions.
Now, assume that the robot chooses a longer route through corridor B instead of a shorter path through corridor A. When asked why it did not take the more direct route through corridor A, the robot could provide different (monolithic) explanations. For instance, “Corridor A likely has high foot traffic at this time of day, which would impede movement and potentially require frequent stops”, or “There is an ongoing maintenance operation in corridor A that creates an obstruction risk and would likely result in a significant detour mid-route.” Both explanations may increase our understanding of why the robot avoided corridor A, but they focus on different aspects of the robot’s belief state and might have different explanatory power.
For the model reconciliation case, consider a human supervisor who expected the robot to take corridor A because they believe the corridor should be clear at this time. This discrepancy arises because the supervisor and the robot have different beliefs about the current state of the building. Then, a model reconciling explanation would need to address these specific belief differences and help the supervisor understand why the robot’s decision was reasonable given its information.
This example illustrates how uncertainty naturally arises in practical scenarios and how multiple potential explanations must be compared to identify those with the highest explanatory power.
@@ 3.1 Propositional Logic
Consider the knowledge base $\mathrm{KB}=\{p,\lnot p\vee q,\lnot p\vee r\}$ built up from $\mathcal{V}=\{p,q,r\}$. Notice that $\mathrm{KB}\models q$. Then, $\epsilon=\{p,\lnot p\vee q\}$ is a monolithic explanation for $q$ from $\mathrm{KB}$.
@@ 3.1.1 Duality of Minimal Unsatisfiable and Minimal Corrections Sets
Partial MUSes in an inconsistent knowledge base $\mathrm{KB}$ appear when a subset of formulae is set as hard, that is, formulae that must always be satisfied in a solution. Conversely, soft formulae may not always be satisfied. Given a formula $\varphi$, we will write $\varphi^{*}$ with $\in\{s,h\}$ to denote it as soft and hard, respectively.
MUSes and MCSes are related by the concept of hitting set:
Definition 5 (Hitting Set).
Given a collection $\Gamma$ of sets from a universe $U$, a hitting set for $\Gamma$ is a set $H\subseteq U$ such that $\forall S\in\Gamma$, $H\cap S\neq\emptyset$. A hitting set $H$ is minimal if $\nexists H^{\prime}\subset H$ such that $H^{\prime}$ is a hitting set.
The relationship between MUSes and MCSes is discussed by ? (?) and ? (?), and it was first presented by ? (?), where MUSes and MCSes are referred to as (minimal) conflicts and diagnoses, respectively.
@@
Let $\mathrm{KB}=\{p,\lnot p\vee q,\lnot p\vee r\}$ and $\epsilon=\{p,\lnot p\vee q\}$ from Example 1. Notice how $\mathcal{M}=\{p,\lnot p\vee q,\lnot q\}$ is an MUS of $\mathrm{KB}\cup\{\lnot q\}$. Then, it is easy to see that $\epsilon$ is a partial MUS of $\mathcal{M}$.
@@ 3.2 Modeling Uncertainty in Propositional Logic
Let $\Omega$ be the set of possible worlds and $P$ a probability distribution over $\Omega$. The degree of belief in a formula $\varphi\in\mathcal{L}$ is $P(\varphi)=\underset{\omega\models\varphi}{{{\sum}}}P(\omega)$. [^3] Note that, with a slight abuse of notation, we use $\omega\models\varphi$ to denote that $\varphi$ is true (e.g., satisfied) in world $\omega$.
@@
$\mathcal{B}=\{(\phi_{1},w_{1}),\ldots,(\phi_{n},w_{n})\}$ (1)
where each formula $\phi_{i}\in\mathcal{L}$ is associated with a weight $w_{i}\in\mathbb{R}^{+}$. [^4] We assume, without loss of generality, that all weights are non-negative because a formula with a negative weight $w$ can be replaced by its negation with weight w-w.
Intuitively, the weights serve as meta-information and reflect the certainty about the truth of the corresponding formulae—the higher the weight, the more certain the formula is. In that sense, formulae with higher weights are prioritized for satisfaction, effectively capturing the certainty of the particular formulae. This mechanism is especially useful for handling inconsistency and non-monotonic reasoning patterns, thus capturing a broader spectrum of problems. [^5] For example, the notion of inconsistency is relaxed as follows: Given two formulae $\phi$ and $\lnot\phi$ that contradict each other (e.g., $\{\phi,\neg\phi\}$ is inconsistent), if $P(\phi)=0.9$, then from the axioms of probability we have that $P(\lnot\phi)=0.1$. This then means that the worlds where $\lnot\phi$ is true are more unlikely than the worlds where $\phi$ is true, but not impossible. Further, we will denote with $\mathcal{B}^{\downarrow w}$ the classical projection of $\mathcal{B}$, that is, $\mathcal{B}^{\downarrow w}=\{\phi_{i}\>|\>(\phi_{i},w_{i})\in\mathcal{B}\}$.
@@
$\forall\omega\in\Omega,\;P_{\mathcal{B}}(\omega)=\frac{1}{Z}\exp\left(\>\overset{n}{\underset{i=1}{{{\sum}}}}w_{i}\cdot\mathbb{I}(\omega,\phi_{i})\right)$ (2)
where $\mathbb{I}(\omega,\phi)=1$ if $\omega\models\phi$ and 0 otherwise, and $Z=\underset{\omega\in\Omega}{\sum}\exp\left(\>\overset{n}{\underset{i=1}{\sum}}w_{i}\cdot\mathbb{I}(\omega,\phi_{i})\right)$ is the normalization factor.
@@
Finally, the weighted formulae in a belief base $\mathcal{B}$ can be viewed as soft constraints in the sense described in Section 3.1.1. In contrast, hard constraints can be imposed as formulae with “infinite” weights. [^6] In practice, infinite weights can be replaced with $\overset{n}{\underset{i=1}{\sum}}w_{i}+1$.
@@ 3.3 The Model Reconciliation Problem
Now, building upon the MRP foundation, we introduced its logic-based variant (L-MRP), where the models of the agent and the human user are represented as logical knowledge bases (?). [^7] We have also introduced extensions that are complementary to the work presented in this paper (?, ?). As a model reconciling explanation must take into account both the knowledge base $\mathrm{KB}_{\alpha}$ of the agent providing an explanation as well as the knowledge base $\mathrm{KB}_{h}$ of the human receiving the explanation, it is defined slightly differently compared to monolithic explanations defined by Definition 1:
@@
Let $\mathrm{KB}_{\alpha}=\{a,\lnot a\vee b,\lnot a\vee c\}$ and $\mathrm{KB}_{h}=\{\lnot a,\lnot a\vee b\}$ be the knowledge bases of an agent and a human user, respectively, where $\mathrm{KB}_{\alpha}\models b$ and $\mathrm{KB}_{h}\not\models b$. A model reconciling explanation is then $\mathcal{E}=\langle\{a\},\{\lnot a\}\rangle$, where $(\mathrm{KB}_{h}\cup\{a\})\setminus\{\lnot a\}=\{a,\lnot a\vee b\}\models b$.
4 Motivating Application: Office Robot Delivery
To demonstrate our ideas, let us revisit the service robot scenario described earlier. Consider a robot operating in an office building with the task of delivering a package from its current location (Room 1) to a destination (Room 2). The robot can navigate through two possible corridors: corridor A, which is shorter but often crowded, or corridor B, which is longer but typically less crowded. Figure 2 illustrates a simplified abstraction of this problem in the form of a grid world.
The robot’s action dynamics include two primary actions: move and deliver. The move action allows the robot to navigate between rooms through a specified corridor, subject to preconditions (the robot must be at the origin location) and probabilistic effects that depend on the corridor’s crowding status, i.e., when attempting to move through a crowded corridor, the robot has a higher probability of failing to move and remaining in its current location. The deliver action can only be performed when the robot is at the destination room and results in the package being delivered. The complete problem specification and encoding is provided in the appendix.
We encode this probabilistic planning scenario as a belief base $\mathcal{B}$ containing weighted formulae (with subscripts denoting time steps) that represent the robot’s beliefs about the environment states and actions. For instance, the initial states and the action dynamics for the move action (for $t=0$) are:
Initial state beliefs:
$(\texttt{robot}\textrm{-}\texttt{at}(room1)_{0},\infty),$
$(\neg\texttt{package}\textrm{-}\texttt{delivered}_{0},\infty),$
$(\texttt{crowded}(A),3),$
$(\texttt{crowded}(B),0.5),$
Move preconditions:
$(\texttt{move}(room1,room2,A)_{0}\rightarrow\texttt{robot}\textrm{-}\texttt{at}(room1)_{0},\infty),$
$(\texttt{move}(room1,room2,B)_{0}\rightarrow\texttt{robot}\textrm{-}\texttt{at}(room1)_{0},\infty),$
Move effects:
$(\texttt{move}(room1,room2,A)_{0}\wedge\texttt{crowded}(A)\rightarrow\texttt{robot}\textrm{-}\texttt{at}(room2)_{1},0.5),$
$(\texttt{move}(room1,room2,A)_{0}\wedge\texttt{crowded}(A)\rightarrow\texttt{robot}\textrm{-}\texttt{at}(room1)_{1},3).$
Figure 2: Grid world illustration of the package delivery problem. The robot has to deliver the package from Room 1 to Room 2 using either Corridor A or Corridor B.
In this belief base, the weights encode the robot’s degrees of belief. For instance, the weight of 3 for $crowded(A)$ represents a strong belief that corridor A is crowded, while the weight of 0.5 for $crowded(B)$ indicates a weaker belief that corridor B is crowded. Similarly, the probabilistic action effects encode that when a corridor is crowded, it is less likely that the robot will move through it to the next room. Given its belief base $\mathcal{B}$, the robot computes that the optimal plan is $\pi=\langle\texttt{move}(room1,room2,B)_{0},\texttt{deliver}_{1}\rangle$, with probability $P(\pi)=0.6$.
Now, suppose a human supervisor observes the robot’s choice and asks: “Why didn’t you take corridor A instead?” This query can be formalized as the explanandum $\varphi=\neg\texttt{move}(room1,room2,A)$, essentially encoding “not moving from Room 1 to Room 2 through corridor A”, and allowing the robot to generate an explanation of why that is the case. In the following section, we will revisit this scenario and show how our methods can be used to generate probabilistic explanations that answer the query.